OpenStudy (anonymous):

Which number is between 0.5 and 5/8 ? A.1/4 B.8/28 C.15/28 D.19/28 Which number is between 1/7 and 0.2 ? A.9/35 B.6/32 C.2/9 D.1/9

OpenStudy (anonymous):

A and A

OpenStudy (nirmalnema):

0.5 and 0.6 = (5/8) 8/28=4/7=5.7approx

OpenStudy (nirmalnema):

1/7=0.1 0.1 and 0.2 =6/32=0.18 som// and 1/9 = 0.11111 greater than 0.1